solve the elimination method x + 8y =14 -x +8y =2
Answer by swincher4391(1107)

x + 8y = 14
-x + 8y = 2
ADD BOTH EQUATIONS TOGETHER
------------
0 + 16y =16
y = 1
-------
Plug y =1 back into either original equation
x + 8(1) = 14
x + 8 = 14
x= 14 -8
x =6
So x = 6, y =1.